Calculate the normal selling price for the top of the line Hammer Time collectors set given the following information.

Direct materials $37/unit

Direct labor $41/unit

Factory Overhead $262,000

Selling & Admin expenses $444,000

Total assets $2,000,000

Base it on 10,000 units

Desired rate of return 17%

The calculated sales price of the collectors set is?

$121.91/unit

$166.31/unit

$173.86/unit

$182.60/unit
Answer #1

Total Variable cost per unit

Direct Material = (37*10,000) = $370,000

Direct Labor = (41 *10,000) = $410,000

Fixed Overhead = $262,000

Selling & Admin expenses = $444,000

Total Cost = $1,486,000

Add: Desired Profit ($2,000,000 * 17%) = $340,000

Total sales = $1,826,000

Selling Price per unit = $1,826,000 / 10,000

Selling Price per unit = $182.60

Option 'D' is correct